You have $250,000 to invest in a stock portfolio. Your choices are Stock H, with an expected return of 13.4 percent, and Stock L, with an expected return of 10.2 percent. If your goal is to create a portfolio with an expected return of 11.3 percent, how much money will you invest in Stock H? In Stock L?
Answer #1

Let investment in H=$x

Hence investment in L=(250,000-x)

Portfolio return=Respective return*Respective weight

11.3=(x/250,000*13.4)+(250,000-x)/250,000*10.2

(11.3*250,000)=13.4x+(250,000-x)*10.2

2,825,000=13.4x+2,550,000-10.2x

Hence x=(2,825,000-2,550,000)/(13.4-10.2)

=$85937.5=investment in H

Hence investment in L=(250,000-x)

=$164062.5